# Analysis of a finite element method for second order uniformly elliptic PDEs in non-divergence form

### v2

@@ 1.1. Literature review
There are several papers [1, 3, 6, 8, 12, 15, 19, 27, 29, 33, 32, 37] focusing on convergence to the viscosity solutions. In the Barles–Souganidis framework [1], monotonicity serves as a fundamental principle underpinning numerical schemes to ensure convergence to viscosity solutions of fully nonlinear partial differential equations. This concept is also applicable to the monotone finite difference methods for the HJB equation [3, 19]. A monotone finite element like scheme for the HJB equation was proposed in [6]. In [32], Nochetto and Zhang introduced a two-scale method based on the integro-differential formulation of (1), where a discrete Alexandroff–Bakelman–Pucci estimate was established. This two-scale method [32] was extended to the Isaacs equation [37]. However, up to our knowledge, all these numerical methods for convergence to the viscosity are of low order accuracy. In addition, most of them have stencils which increase under mesh refinement, such that the corresponding linear system may not be sparse.
@@
    AA can be discontinuous but $\gamma A$ dominated by $I_{d}$ where $\gamma$ is a positive weight function. In [39], Smears and Süli introduced and analyzed a Discontinuous Galerkin (DG) method for (1). The main assumptions are the coefficients $A$ is dominated by $I_{d}$ after being multiplied by a positive weight function $\gamma$ (Cordès condition) and the domain $\Omega$ is convex. The stability and optimal rate of convergence in discrete $H^{2}$-norm were proven in [39]. Following this direction, there are several papers [2, 21, 43]. The technique in [39] has been extended to the HJB equation (2) in [40, 22, 43]. It is not clear how to prove optimal rate of convergence in discrete $W^{2,p}$-norm ($p\neq 2$) for this kind of numerical methods.
@@
    AA is globally continuous. In [14], Feng and Hennings and Neilan introduced and analyzed a $C^{0}$-conforming DG method for (1). The main assumptions are $A\in C^{0}(\overline{\Omega})$ and the domain $\Omega$ has $C^{1,1}$ boundary. They followed the approach in [24, Chapter $9$] to prove that their numerical method could achieve stability and optimal rate of convergence in discrete $W^{2,p}$-norm for any $1<p<+\infty$. However, this numerical method and its variant - a DG method in [17] can not deal with discontinuous coefficients $A$, since they have the jump of $A\nabla u_{h}$ along mesh interfaces ($u_{h}$ is the numerical solution). Therefore, the numerical methods in [14, 17] can not be applied for the HJB equation (2).
@@ 1.2. Main contributions
We introduce the finite element methods (14) and (16) for the linear elliptic PDE in non-divergence form (1) and the HJB equation (2), respectively. It is easy to see that the FEM (16) becomes (14), if the index set $\Lambda$ has a single element. Therefore, we can claim that essentially, there is only one FEM for (1) and (2). We would like to point out that the discrete gradient operator $\overline{\nabla}_{h}$ (see (11)) used in (14) and (16) has been introduced in [2, 11, 18].
@@
Here $\epsilon_{1}$ and $\epsilon_{2}$ are two positive numbers up to the domain $\Omega$ (see Lemma 2.1).
@@ 1.2.1. Contributions for the linear elliptic PDE in non-divergence form (1)
When the coefficient matrix $A\in C^{0}(\overline{\Omega})$, Theorem 3.4 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ and optimal convergence with respect to $W_{h}^{2,p}$-norm (see (7)) of the numerical solution of the FEM (14). The well-posedness of (1) provided by [24, Theorem $9.15$] and optimal convergence of the FEM in [14] require $\partial\Omega$ to be $C^{1,1}$. In contrast, Theorem 3.4 is valid on Lipschitz polyhedral domains (see (5) for the detailed description). In fact, with the help of the well-posedness of strong solution stated in Theorem 3.4, the $W_{h}^{2,p}$-norm stability of the FEM in [14] can be immediately extended from domains with $C^{1,1}$ boundary to Lipschitz polyhedral domains with $p$ described in (5b). Furthermore, in contrast to the proof of numerical stability presented in [14], which rests upon a specialized duality argument (see the proof of [14, Lemma $3.5$]) contingent on the well-posedness of (1), the present work establishes numerical stability (Theorem 3.3) by invoking the discrete compactness property (Lemma 3.8) and the uniqueness of the strong solution of (1). Thus, the analysis of this paper is quite different from that of [14].
@@
    where the range of $p$ is described in (5a). The most difficult part of the proof of Theorem 3.2 is to obtain the above estimate for $1<p\leq\frac{3}{2}$ if $\Omega$ is a convex polyhedra in $\mathbb{R}^{3}$. It is because the Alexandroff–Bakelman–Pucci estimate can not be applied directly due to the fact $W^{2,p}(\Omega)$ can not be embedded into $C^{0}(\overline{\Omega})$. To overcome this difficulty, we provide Lemma 3.1 which shows the uniqueness of the strong solution of (1) in $W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$. Then Theorem 3.2 basically follows [34, Theorem $1.1$] by a proof by contraction.
@@
    Here, $V_{h}$ is the finite element space introduced in Section 2.1, $\mathcal{L}_{\tilde{A},h}$ is the Riesz representation (see (40)) of the FEM (14) with $\boldsymbol{b}=\boldsymbol{0}$ and $c=0$. The proof of Lemma 3.7 mimics that of [14, Lemma $3.4$]. However, due to the construction of the discrete gradient operator (11), the proof of Lemma 3.7 is much more technical.
  • Step $3$. To erase the term $\|w_{h}\|_{W^{1,p}(\Omega)}$ on the right hand side of (42), we prove by contradiction with the help of the discrete compactness - Lemma 3.8 and the uniqueness of strong solution of (1) by Theorem 3.2. Then we obtain Theorem 3.2 which the stability of the FEM (14):
@@
    We would like to point out that to obtain the same result, the authors of [14] used a specialized duality argument (see the proof of [14, Lemma $3.5$]), which relies on the well-posedness of strong solution of (1).
  • Step $4$. Roughly speaking, we apply the discrete compactness - Lemma 3.8 with Theorem 3.2 to prove the limit of numerical solution of the FEM (14) is a solution of (1). Then by the global $W^{2,p}$ estimate (26), the well-posedness of strong solution in $W^{2,p}(\Omega)$ of (1) is proven for the range of $p$ described in (5b). Finally by some density argument, we extend $p$ to the range described in (5a). Optimal convergence of the numerical solution of the FEM (14) is straightforward.
@@ 1.2.2. Contributions for the elliptic HJB equation (2)
For the elliptic HJB equation (2), the main theoretical results are Theorem 4.1 and Theorem 4.3. Theorem 4.1 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ of the elliptic HJB equation (2), if the coefficients satisfy the condition (83). In parallel, Theorem 4.3 provides optimal convergence of the FEM (16) in $W_{h}^{2,p}$-norm, if the coefficients satisfy the condition (93). When the domain is convex and $p=2$, the condition (83) is identical to the Cordès condition used in [40] but the condition (93) is more restrictive. Please see Remark 4.2 and Remark 4.4 for the special case $\boldsymbol{b}^{\alpha}=\boldsymbol{0}$ and $c^{\alpha}=0$ for any $\alpha\in\Lambda$.
We would like to emphasize that [40] assumes $A^{\alpha}\in C^{0}(\overline{\Omega}\times\Lambda)$ and the index set $\Lambda$ is a compact metric space, while [23, Proposition $4.3$] assumes that there are $\beta\in(0,1]$ and $0<C<+\infty$ such that $\|A^{\alpha}\|_{C^{0,\beta}(\overline{\Omega})}<C$ for any $\alpha\in\Lambda$ ([23, Proposition $4.3$] assumes $\boldsymbol{b}^{\alpha}=\boldsymbol{0}$, $c^{\alpha}=0$, and $f^{\alpha}=f$). So, if the index set $\Lambda$ has a single element, the elliptic HJB equation (2) will not become the linear elliptic PDE in non-divergence form (1) with discontinuous matrix coefficient $A$. In contrast, the assumption (6) allows discontinuity of $A^{\alpha}$, $\boldsymbol{b}^{\alpha}$ and $c^{\alpha}$. By Lemma 4.3, we can conclude that the assumptions used by [40] and [23, Proposition $4.3$] imply that the assumption (6) holds.
If the index set $\Lambda$ is countable, then for any $w\in W^{2,1}(\Omega)$, the function
$\sup_{\alpha\in\Lambda}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}]$
is a measurable function on $\Omega$. However, it is not obviously true if $\Lambda$ is uncountable. Lemma 4.1 shows that the assumption (6) implies (see (78))
$\sup_{\alpha\in\Lambda}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}]=\sup_{\tilde{\alpha}\in\tilde{\Lambda}}[A^{\tilde{\alpha}}:D^{2}w+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla w+c^{\tilde{\alpha}}w-f^{\tilde{\alpha}}$],
where $\tilde{\Lambda}$ is a countable subset of $\Lambda$. Thus this issue is solved with the assumption (6) which is more general than assumptions in [40] and [23, Proposition $4.3$].
@@ 2.1. Mesh and space notation
Associated with $D\subset\Omega$, we define a semi-norm on $W_{h}^{2,p}(D)$ for $1<p<\infty$:
@@
$\|v\|_{W_{h}^{2,p}(\Omega)}=\|D_{h}^{2}v\|_{L^{p}(\Omega)}+\left(\Sigma_{F\in\mathcal{F}_{h}^{I}}h_{F}^{1-p}\|[\nabla v]\|_{L^{p}(F)}^{p}\right)^{\frac{1}{p}}.$
@@
We would like to point out that the discrete gradient operator (11) has been introduced in [2, 11, 18].
@@ 2.2.1. FEM for linear elliptic PDE in non-divergence form
$=$ $\int_{\Omega}\gamma(\tilde{A}:\overline{\nabla}_{h}(\nabla u_{h})+\tilde{\boldsymbol{b}}\cdot\nabla u_{h}+\tilde{c}u_{h})v_{h}d\boldsymbol{x}=\int_{\Omega}\gamma fv_{h}d\boldsymbol{x}=\int_{\Omega}\tilde{f}v_{h}d\boldsymbol{x},\qquad\forall v_{h}\in V_{h}.$
@@ 2.3. Preliminary results
The proof of existence of a unique solution and the estimate (17) are due to [7, Proposition $2.1$] ($d=2$) and [20, ($1.6$)] ($d=3$). The proof of existence of a unique solution and the estimate (18) are due to [28, Theorem $0.5$] (non-convex and $d=2$) and [20, ($1.5$)] (convex). The estimate (19) is due to [39, Lemma $1$]. ∎
According to [9, Theorem $3$ and Theorem $4$] (see [42, Lemma $7.3$] as well), we have the following Lemma 2.2 since we assume the mesh $\mathcal{T}_{h}$ is quasi-uniform. We would like to point out that the proof of [9, Theorem $4$] can be easily extended to obtain (20b) in three dimensional domains.
@@
According to [13, Lemma $4.3$], we have the following Lemma 2.3.
@@
For any $\overline{v}_{h}\in\overline{V}_{h}$, we denote by $\chi_{h}\in H^{1}(\Omega)\cap\overline{V}_{h}$ the standard $H^{1}$-conforming interpolation of $\overline{v}_{h}$. Then for any $q\in[1,+\infty]$, we have

### v3

@@ 1.1. Literature review
There are several papers [1, 3, 7, 9, 13, 16, 20, 28, 30, 36, 35, 40] focusing on convergence to the viscosity solutions. In the Barles–Souganidis framework [1], monotonicity serves as a fundamental principle underpinning numerical schemes to ensure convergence to viscosity solutions of fully nonlinear partial differential equations. This concept is also applicable to the monotone finite difference methods for the HJB equation [3, 20]. A monotone finite element like scheme for the HJB equation was proposed in [7]. In [35], Nochetto and Zhang introduced a two-scale method based on the integro-differential formulation of (1), where a discrete Alexandroff–Bakelman–Pucci estimate was established. This two-scale method [35] was extended to the Isaacs equation [40]. However, up to our knowledge, all these numerical methods converge to the viscosity with low order accuracy. In addition, most of them have stencils which increase under mesh refinement, such that the corresponding linear system may not be sparse.
@@
    AA can be discontinuous but $\gamma A$ dominated by $I_{d}$ where $\gamma$ is a positive weight function. In [42], Smears and Süli introduced and analyzed a Discontinuous Galerkin (DG) method for (1). The main assumptions are the coefficients $A$ is dominated by $I_{d}$ after being multiplied by a positive weight function $\gamma$ (Cordes condition) and the domain $\Omega$ is convex. The stability and optimal rate of convergence in discrete $H^{2}$-norm were proven in [42]. Following this direction, there are several papers [2, 22, 46]. The technique in [42] has been extended to the HJB equation (2) in [43, 23, 46]. It is not clear how to prove optimal rate of convergence in discrete $W^{2,p}$-norm ($p\neq 2$) for this kind of numerical method.
@@
    AA is globally continuous. In [15], Feng, Hennings and Neilan introduced and analyzed a $C^{0}$-conforming DG method for (1). The main assumptions are $A\in C^{0}(\overline{\Omega})$ and the domain $\Omega$ has $C^{1,1}$ boundary. They followed the approach in [25, Chapter $9$] to prove that their numerical method achieves stability and optimal rate of convergence in discrete $W^{2,p}$-norm for any $1<p<+\infty$. However, this numerical method and its variant - a DG method in [18] can not deal with discontinuous coefficients $A$, since they have the jump of $A\nabla u_{h}$ along mesh interfaces ($u_{h}$ is the numerical solution). Therefore, the numerical methods in [15, 18] can not be applied for the HJB equation (2).
@@ 1.2. Main contributions
We introduce the finite element methods (14) and (16) for the linear elliptic PDE in non-divergence form (1) and the HJB equation (2), respectively. It is easy to see that the FEM (16) becomes (14), if the index set $\Lambda$ has a single element. Therefore, we can claim that essentially, there is only one FEM for (1) and (2). We would like to point out that the discrete gradient operator $\overline{\nabla}_{h}$ (see (11)) used in (14) and (16) has been introduced in [31], and the finite element method (14) is similar to the one in [31]. [34] provides optimal convergence in discrete $H^{2}$-norm for the method in [31] for (1) with $A\in C^{0}(\overline{\Omega})$ on domains with $C^{1,1}$ boundary. The technique in [34] is similar to those in [15, 18]. It is not clear how to generalize the analysis in [34] for (2).
@@
Here $\epsilon_{1}$ and $\epsilon_{2}$ are two positive numbers which depend on the domain $\Omega$ (see Lemma 2.1).
@@ 1.2.1. Contributions for the linear elliptic PDE in non-divergence form (1)
When the coefficient matrix $A\in C^{0}(\overline{\Omega})$, Theorem 3.3 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ and optimal convergence with respect to $W_{h}^{2,p}$-norm (see (7)) of the numerical solution of the FEM (14). The well-posedness of (1) provided by [25, Theorem $9.15$] and optimal convergence of the FEM in [15] require $\partial\Omega$ to be $C^{1,1}$. In contrast, Theorem 3.3 is valid on Lipschitz polyhedral domains (see (5) for the detailed description). In fact, with the help of the well-posedness of strong solution stated in Theorem 3.3, the $W_{h}^{2,p}$-norm stability of the FEM in [15] can be immediately extended from domains with $C^{1,1}$ boundary to Lipschitz polyhedral domains with $p$ described in (5b). Furthermore, in contrast to the proof of numerical stability presented in [15], which rests upon a specialized duality argument (see the proof of [15, Lemma $3.5$]) contingent on the well-posedness of (1), the present work establishes numerical stability (Theorem 3.2) by invoking the discrete compactness property (Lemma 3.8) and the uniqueness of the strong solution of (1). Thus, the analysis of this paper is quite different from that of [15, 18, 34].
@@
    where the range of $p$ is described in (5a). The most difficult part of the proof of Theorem 3.1 is to obtain the above estimate for $1<p\leq\frac{3}{2}$ if $\Omega$ is a convex polyhedra in $\mathbb{R}^{3}$. It is because the Alexandroff–Bakelman–Pucci estimate can not be applied directly due to the fact $W^{2,p}(\Omega)$ can not be embedded into $C^{0}(\overline{\Omega})$. To overcome this difficulty, we provide Lemma 3.1 which shows the uniqueness of the strong solution of (1) in $W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$. Then Theorem 3.1 basically follows [37, Theorem $1.1$] by a proof by contraction. Though it is inspired by that of [25, Lemma $9.16$], the proof of Lemma 3.1 avoids using the common technique of “flattening” $\partial\Omega$, which is critical in [25, Lemma $9.16$]. Thus Lemma 3.1 is valid on convex polyhedra in $\mathbb{R}^{3}$ (It is trivial to obtain the uniqueness on two dimensional Lipschitz polygon).
@@
    Here, $V_{h}$ is the finite element space introduced in Section 2.1, $\mathcal{L}_{\tilde{A},h}$ is the Riesz representation (see (47)) of the FEM (14) with $\boldsymbol{b}=\boldsymbol{0}$ and $c=0$. The proof of Lemma 3.7 mimics that of [15, Lemma $3.4$]. However, due to the construction of the discrete gradient operator (11), the proof of Lemma 3.7 is much more technical.
  • Step $3$. To erase the term $\|w_{h}\|_{W^{1,p}(\Omega)}$ on the right hand side of (49), we prove by contradiction with the help of the discrete compactness - Lemma 3.8 (can be considered as an analogue of [4, Theorem $5.2$] for the discrete $W^{2,p}$-norm) and the uniqueness of strong solution of (1) by Theorem 3.1. Then we obtain Theorem 3.1 which the stability of the FEM (14):
@@
    We would like to point out that to obtain the same result, the authors of [15] used a specialized duality argument (see the proof of [15, Lemma $3.5$]), which relies on the well-posedness of strong solution of (1). Actually, our approach in this step is a discrete analogue of [25, Lemma $9.17$]. Furthermore, to prove Lemma 3.8, we provide Lemma 2.2 which shows for any $p\in[1,+\infty]$,
    $\|v_{h}\|_{W^{1,p}(\Omega)}\leq C\|v_{h}\|_{W_{h}^{2,p}(\Omega)},\qquad\forall v_{h}\in V_{h}.$
    Here the constant $C$ is independent of $p\in[1,+\infty]$.
  • Step $4$. Roughly speaking, we apply the discrete compactness - Lemma 3.8 with Theorem 3.1 to prove the limit of numerical solution of the FEM (14) is a solution of (1). Then by the global $W^{2,p}$ estimate (32), the well-posedness of a strong solution to (1) is proven for the range of $p$ described in (5b). Finally by some density argument, we extend $p$ to the range described in (5a). Optimal convergence of the numerical solution of the FEM (14) is straightforward.
@@ 1.2.2. Contributions for the elliptic HJB equation (2)
For the elliptic HJB equation (2), the main theoretical results are Theorem 4.1 and Theorem 4.3. Theorem 4.1 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ of the elliptic HJB equation (2), if the coefficients satisfy the condition (90). In parallel, Theorem 4.3 provides optimal convergence of the FEM (16) in $W_{h}^{2,p}$-norm, if the coefficients satisfy the condition (100). When the domain is convex and $p=2$, the condition (90) is identical to the Cordes condition used in [43] but the condition (100) is more restrictive. Please see Remark 4.2 and Remark 4.4 for the special case $\boldsymbol{b}^{\alpha}=\boldsymbol{0}$ and $c^{\alpha}=0$ for any $\alpha\in\Lambda$.
We would like to emphasize that [43] assumes $A^{\alpha}\in C^{0}(\overline{\Omega}\times\Lambda)$ and the index set $\Lambda$ is a compact metric space, while [24, Proposition $4.3$] assumes that there are $\beta\in(0,1]$ and $0<C<+\infty$ such that $\|A^{\alpha}\|_{C^{0,\beta}(\overline{\Omega})}<C$ for any $\alpha\in\Lambda$ ([24, Proposition $4.3$] assumes $\boldsymbol{b}^{\alpha}=\boldsymbol{0}$, $c^{\alpha}=0$, and $f^{\alpha}=f$). So, if the index set $\Lambda$ has a single element, the elliptic HJB equation (2) will not become the linear elliptic PDE in non-divergence form (1) with discontinuous matrix coefficient $A$. In contrast, the assumption (6) allows discontinuity of $A^{\alpha}$, $\boldsymbol{b}^{\alpha}$ and $c^{\alpha}$. By Lemma 4.3, we can conclude that the assumptions used by [43] and [24, Proposition $4.3$] imply that the assumption (6) holds.
We notice that for any $\alpha\in\Lambda$, $\boldsymbol{b}^{\alpha}$ and $c^{\alpha}$ may be discontinuous. To have one numerical method for the linear elliptic PDE in non-divergence form (1) and the HJB equation (2), the positive weight function $\gamma^{\alpha}$ in (15) can not depend on $\boldsymbol{b}^{\alpha}$ and $c^{\alpha}$. Otherwise the analysis for the FEM (14) for (1) will not be valid for $A\in C^{0}(\Omega)$. Furthermore, we don’t want to include the parameter $\lambda$ used in the similar weight function in [43] which might not be known beforehand. These restrictions make the analysis of the FEM (16) for (2) more difficult than that in [43].
@@ 2.1. Mesh and space notation
Associated with $D\subset\Omega$, we define a semi-norm on $W_{h}^{2,p}(D)$ for $p\in[1,+\infty]$:
@@
$\|v\|_{W_{h}^{2,p}(\Omega)}=\|D_{h}^{2}v\|_{L^{p}(\Omega)}+\left(\Sigma_{F\in\mathcal{F}_{h}^{I}}h_{F}^{1-p}\|[\nabla v]\|_{L^{p}(F)}^{p}\right)^{\frac{1}{p}},$
$\|v\|_{W_{h}^{2,+\infty}(\Omega)}=\|D_{h}^{2}v\|_{L^{\infty}(\Omega)}+\max_{F\in\mathcal{F}_{h}^{I}}h_{F}^{-1}\|[\nabla v]\|_{L^{\infty}(F)}.$
@@
We would like to point out that the discrete gradient operator (11) has been introduced in [31]. It has been used in [2, 12, 19] as well.
@@ 2.2.1. FEM for linear elliptic PDE in non-divergence form
$=$ $\int_{\Omega}\gamma(A:\overline{\nabla}_{h}(\nabla u_{h})+\boldsymbol{b}\cdot\nabla u_{h}+cu_{h})v_{h}d\boldsymbol{x}=\int_{\Omega}\gamma fv_{h}d\boldsymbol{x}=\int_{\Omega}\tilde{f}v_{h}d\boldsymbol{x},\qquad\forall v_{h}\in V_{h}.$
@@ 2.3. Preliminary results
The proof of existence of a unique solution and the estimate (17) are due to [8, Proposition $2.1$] ($d=2$) and [21, ($1.6$)] ($d=3$). The proof of existence of a unique solution and the estimate (18) are due to [29, Theorem $0.5$] (non-convex and $d=2$) and [21, ($1.5$)] (convex). The estimate (19) is due to [42, Lemma $1$]. ∎
According to [10, Theorem $3$ and Theorem $4$] (see [45, Lemma $7.3$] as well), we have the following Lemma 2.2 since we assume the mesh $\mathcal{T}_{h}$ is quasi-uniform. We would like to point out that the proof of [10, Theorem $4$] can be easily extended to obtain (20b) in three dimensional domains.
@@
According to [14, Lemma $4.3$], we have the following Lemma 2.3.
@@
For any $\overline{v}_{h}\in\overline{V}_{h}$, we denote by $\chi_{h}\in H^{1}(\Omega)\cap\overline{V}_{h}$ the standard $H^{1}$-conforming averaging of $\overline{v}_{h}$. Then there is a constant $C$ such that for any $q\in[1,+\infty]$, we have
@@
Here the domain $\Omega$ can be any Lipschitz polyhedra in $\mathbb{R}^{d}$ ($d=2,3$).
Lemma 2.4.
There is a constant $C$ such that for any $q\in[1,+\infty]$, we have
$\|\nabla v_{h}\|_{L^{q}(\Omega)}\leq C\|v_{h}\|_{W_{h}^{2,q}(\Omega)},\qquad\forall v_{h}\in V_{h}.$
Here the domain $\Omega$ can be any Lipschitz polyhedra in $\mathbb{R}^{d}$ ($d=2,3$).
Proof.
We choose $v_{h}\in V_{h}$ and $q\in[1,+\infty]$ arbitrarily. We denote by $C$ a constant independent of $v_{h}$ and $q\in[1,+\infty]$. According to Lemma 2.3, there is $\boldsymbol{\psi}_{h}\in[H^{1}(\Omega)\cap P_{r}(\mathcal{T}_{h})]^{d}$ such that
(21) $\|\boldsymbol{\psi}_{h}-\nabla v_{h}\|_{L^{q}(\Omega)}\leq Ch\left(\Sigma_{F\in\mathcal{F}_{h}^{I}}h_{F}^{1-q}\|[\nabla v_{h}]\|_{L^{q}(F)}^{q}\right)^{\frac{1}{q}}.$
Since the proof for the case $d=2$ is similar but simpler, we assume $d=3$ in the following. Since $v_{h}=0$ on $\partial\Omega$, $\boldsymbol{n}\times(\nabla v_{h})=\boldsymbol{0}$ on $\partial\Omega$. Then by (21), we have
(22) $\|\boldsymbol{n}\times\boldsymbol{\psi}_{h}\|_{L^{q}(\partial\Omega)}=\|\boldsymbol{n}\times(\boldsymbol{\psi}_{h}-\nabla v_{h})\|_{L^{q}(\partial\Omega)}$
$\leq$ $Ch^{1-\frac{1}{q}}\left(\Sigma_{F\in\mathcal{F}_{h}^{I}}h_{F}^{1-q}\|[\nabla v_{h}]\|_{L^{q}(F)}^{q}\right)^{\frac{1}{q}}\leq C\|v_{h}\|_{W_{h}^{2,q}(\Omega)}.$
Let $\{\psi_{h}^{(i)}\}_{i=1}^{3}$ be the collection of all components of $\boldsymbol{\psi}_{h}$. Since the domain $\Omega$ is a polyhedra, we can assume that there is a face $\tilde{F}$ on $\partial\Omega$ which is parallel to both $x_{1}$-axis and $x_{2}$-axis. Then by (22), we have
(23) $\|\psi_{h}^{(1)}\|_{L^{q}(\tilde{F})}+\|\psi_{h}^{(2)}\|_{L^{q}(\tilde{F})}\leq C\|v_{h}\|_{W_{h}^{2,q}(\Omega)}.$
In the following, we will prove
(24) $\|\psi_{h}^{(i)}\|_{L^{q}(\Omega)}\leq C\left(\|\psi_{h}^{(i)}\|_{L^{q}(\tilde{F})}+\|\nabla\psi_{h}^{(i)}\|_{L^{q}(\Omega)}\right),\qquad\forall i\in\{1,2\}.$
In fact, if (24) holds, then (21,23) imply $\|\dfrac{\partial v_{h}}{\partial x_{1}}\|_{L^{q}(\Omega)}+\|\dfrac{\partial v_{h}}{\partial x_{2}}\|_{L^{q}(\Omega)}\leq C\|v_{h}\|_{W_{h}^{2,q}(\Omega)}$. Then by applying the same argument to another face on $\partial\Omega$ which is not parallel to $\tilde{F}$, we will have $\|\nabla v_{h}\|_{L^{q}(\Omega)}\leq C\|v_{h}\|_{W_{h}^{2,q}(\Omega)}$.
Therefore, we only need to prove (24). For the sake of simplicity, we will only prove (24) for $i=1$. We decompose the domain $\Omega$ into finitely many non-overlapping simplexes $\{T_{i}\}_{i=1}^{N}$, such that $\overline{\Omega}=\cup_{i=1}^{N}\overline{T}_{i}$. We assume this decomposition of $\Omega$ is conforming such that it is like a conforming mesh of $\Omega$. We would like to point out that this decomposition of $\Omega$ is not relevant to the mesh $\mathcal{T}_{h}$, such that $N$ is independent of $h$. If we take an element $K\in\mathcal{T}_{h}$ arbitrarily, $K$ may have non-trivial overlapping with two or more simplexes among $\{T_{i}\}_{i=1}^{N}$.
[…]
(25) $\|\psi_{h}^{(1)}\|_{L^{q}(T_{i})}+\|\psi_{h}^{(1)}\|_{L^{q}(F_{i}^{\prime\prime})}\leq C\left(\|\psi_{h}^{(1)}\|_{L^{q}(F_{i}^{\prime})}+\|\nabla\psi_{h}^{(1)}\|_{L^{q}(T_{i})}\right),$
where $F_{i}^{\prime}$ and $F_{i}^{\prime\prime}$ can be any two triangular faces of $T_{i}$. If (25) holds, then (24) is true for $i=1$ because (23) holds and $N$ is independent of $h$.
Therefore, we only need to prove (25). Since $\psi_{h}^{(1)}\in H^{1}(\Omega)\cap P_{r}(\mathcal{T}_{h})$, $\psi_{h}^{(1)}$ is globally Lipschitz on $\overline{\Omega}$. So $\psi_{h}^{(1)}$ is globally Lipschitz on $\overline{T_{i}}$ for any $1\leq i\leq N$. We notice that the Lipschitz constant of $\psi_{h}^{(i)}$ on $\overline{T}_{i}$ is bounded by (independent of $h$) $\|\nabla\psi_{h}^{(i)}\|_{L^{\infty}(T_{i})}$ for any $1\leq i\leq N$. So (25) holds for $q=+\infty$.
Now we are going to prove (25) for $q\in[1,+\infty)$. Via affine transformation, we can assume $T_{i}$ is the simplex determined by the four vertexes $(0,0,0)$, $(1,0,0)$, $(0,1,0)$ and $(0,0,1)$. We further assume $F_{i}^{\prime}$ is the triangular face through $(0,0,0)$, $(1,0,0)$ and $(0,1,0)$, and $F_{i}^{\prime\prime}$ is the triangular face through $(1,0,0)$, $(0,1,0)$ and $(0,0,1)$. Since $\psi_{h}^{(i)}$ is globally Lipschitz on $\overline{T}_{i}$, we have that for any $(x_{1},x_{2},x_{3})\in\overline{T}_{i}$,
(26) $\psi_{h}^{(1)}(x_{1},x_{2},x_{3})=\psi_{h}^{(1)}(x_{1},x_{2},0)+\int_{0}^{x_{3}}\dfrac{\partial\psi_{h}^{(1)}}{\partial x_{3}}(x_{1},x_{2},s)ds.$
Since $\psi_{h}^{(1)}\in H^{1}(\Omega)\cap P_{r}(\mathcal{T}_{h})$, we have that for any $(x_{1},x_{2},0)\in\overline{F}_{i}^{\prime}$, $\dfrac{\partial\psi_{h}^{(1)}}{\partial x_{3}}(x_{1},x_{2},s)$ on the right hand side of (26) is the same as the weak derivative of $\psi_{h}^{(i)}$ with respect to $x_{3}$ on the line segment $l(x_{1},x_{2}):=\{(x_{1},x_{2},s):0<s<1-x_{1}-x_{2}\}$ except at most finitely many points. In fact, these points are intersections between the line segment $l(x_{1},x_{2})$ and interior mesh interfaces in $\mathcal{F}_{h}^{I}$ which are not parallel to $l(x_{1},x_{2})$. Therefore, (26) implies that for any $q\in[1,+\infty)$,
$\|\psi_{h}^{(1)}\|_{L^{q}(F_{i}^{\prime\prime})}\leq\left(\sqrt{3}\int_{F_{i}^{\prime}}|\psi_{h}^{(1)}(x_{1},x_{2},1-x_{1}-x_{2})|^{q}dx_{1}dx_{2}\right)^{\frac{1}{q}}$
$\leq$ $\sqrt{3}\left(\int_{F_{i}^{\prime}}|\psi_{h}^{(1)}(x_{1},x_{2},1-x_{1}-x_{2})|^{q}dx_{1}dx_{2}\right)^{\frac{1}{q}}\leq C\left(\|\psi_{h}^{(1)}\|_{L^{q}(F_{i}^{\prime})}+\|\nabla\psi_{h}^{(1)}\|_{L^{q}(T_{i})}\right),$
$\|\psi_{h}^{(1)}\|_{L^{q}(T_{i})}\leq C\left(\|\psi_{h}^{(1)}\|_{L^{q}(F_{i}^{\prime})}+\|\nabla\psi_{h}^{(1)}\|_{L^{q}(T_{i})}\right).$
Therefore, we have that (25) for $q\in[1,+\infty]$. We can conclude that the proof is complete. ∎